[Brahms-l] Shift report 20040104 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 04 2004 - 08:13:23 EST
------------------------------------------------------------
Date: Sun Jan  4 08:13:17 2004, Shift: 00:00-08:00
Supervisor: hiro, with: 
Log book#15 pages 22 - 22

Shift Summary:
------------------------------------------------------------
1.  Beam for entire night.
    The center of vertex distribution seemes to be off (to the left.)  
2.  Run 9349, 9350, 9351.
3.  1 FERA ADC for silicon detector has missing channels.  It needs to be replaced.
    2. BFS does not move at all.  Need Dana to investigate!!!


------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 4 08:13:35 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 04 2004 - 08:13:52 EST